Problem 1

Implementing an e-commerce website has a projected net economic benefit of $ 40,000/year during the next five years. The recurring costs of this project are $ 10,000/year, and the initial cost of implementing the website is $ 50,000. Assuming that 10 percent is the estimated discount rate, do the following:

  1. Calculate the net present value of these costs and benefits
  2. Calculate the overall return on investment
  3. Perform a break-even analysis and estimate when the actual break-even occurs
  4. Generate a graphical representation in the case that the project breaks even
Answer #1
Year Cash flow PV @ 10% Present value
0 -50000 1 $ -50,000.00
1 30000 0.90909091 $ 27,272.73
2 30000 0.82644628 $ 24,793.39
3 30000 0.7513148 $ 22,539.44
4 30000 0.68301346 $ 20,490.40
5 30000 0.62092132 $ 18,627.64
$ 63,723.60
Alternative method
Year= 0 1 2 3 4 5
Benefit 0 40000 40000 40000 40000 40000
Cost 50000 10000 10000 10000 10000 10000
Net -50000 30000 30000 30000 30000 30000
Disc. Factor 1.10^0=1 1.10^1= 1.10^2= 1.10^3= 1.10^4= 1.10^5=
1 1.1 1.21 1.331 1.4641 1.61051
Disc Cash Flow -50000 27272.7273 24793.38843 22539.444 20490.4037 18627.6397
NPV $    63,723.60
Overall return on investment = IRR 52.80%
Discount rate NPV
0% $ 1,00,000.00
10% $    63,723.60
20% $    39,718.36
30% $    23,067.09
40% $    11,054.92
50% $      2,098.77
60% $    -4,768.37
70% $ -10,161.27
80% $ -14,484.58
90% $ -18,012.87
100% $ -20,937.50
BEP = NPV = 0 at IRR 52.80%

NPV Chart Area $80,000.00 $60,000.00 40,000.00 20,000.00 S- 0% 2088 40% 60% 100% 120%